OpenStudy (anonymous):

write an expression using all six trigonometric functions such that the value of the expression is 3

OpenStudy (freckles):

I would us Pythagorean identities but that isn't the only way: \[\sin^2(x)+\cos^2(x)=1 \\ \tan^2(x)+1=\sec^2(x) \implies \sec^2(x)-\tan^2(x)=1\\ 1+\cot^2(x)=\csc^2(x) \implies \csc^2(x)-\cot^2(x)=1\]

OpenStudy (freckles):

1+1+1=3 see if you can finish from here